OpenStudy (anonymous):

How would you integrate (x-x^2)/(2 cube root x) with a lower limit of -1 and upper limit of -8?

OpenStudy (anonymous):

\[\int\limits_{-8}^{-1} \frac{ x-x ^{2} }{ 2\sqrt[3]{x} } dx\]

OpenStudy (anonymous):